In constructing an s chart for monitoring the variability in the number of fries in a package, the estimate of the standard deviation is s 3, based on samples of size 7. The center line for the chart will be at which of the following? 2.88 0.9594 3.00
Answer #1

Centerline Control Limits X bar and R Charts CLR R d2 X bar and s Charts CL, C4

Center line = c4 * σ = 0.9594 * 3 = 2.88